# Dynamical complex networks for understanding motility induced phase separation in Active Brownian Particles

### v1

@@ Abstract
We investigate the dynamics of active Brownian particles (ABP) through a temporal complex network framework, focusing on degree probability distribution, average path length, and clusterization across Motility-Induced Phase Separation (MIPS) regions. In the single MIPS phase, particle encounters resemble a random graph. Transitioning toward the MIPS frontier introduces a non-power-law distribution, while the phase-separated region exhibits a double power-law distribution. Numerical analysis indicates a shift from random to power-law complex networks based on the system’s MIPS-phase position. In summary, our numerical and theoretical analysis reveals the phase transition of the system between two states and how this transition manifests itself in the complex network, showing an evident change between a random network and a network that tends to have scale-free behavior.
1 Introduction
Active Brownian Particles (ABP) serve as a valuable framework for investigating the non-equilibrium dynamics and statistical properties of self-propelled agents, whether they are living organisms or artificial entities romanczuk2012active ; bechinger2016active . In contrast to passive Brownian particles studied in random diffusive systems, ABP offers a new dimension of complexity due to the inherent particle self-propulsion mechanisms. In particular, these active particles undergo a distinctive phase transition known as Motility-Induced Phase Separation (MIPS) fily2012athermal ; redner2013structure ; stenhammar2014phase ; theers2018clustering ; van2019interrupted . This phase transition is related to the spontaneous separation between dilute (free-particles) and dense phases (cluster formation) stenhammar2014phase ; de2023sequential ; de2021active . This phase separation is a fascinating emergent phenomenon where particles collapse into clusters although natural repulsion forces are present. The exploration of ABP not only deepens our understanding of self-organization and emergent phenomena but also opens up promising avenues for harnessing their unique properties in various applications and technological advancements gompper20202020 and building a bridge between physical models and biological systems.
@@ 1 Introduction
2 Active Brownian Particles (ABP) and motility-phase-separation
@@
$\dot{\mathbf{r}}_{i}$ $=$ $u_{0}\hat{\mathbf{u}}_{i}(t)+\bm{\xi}_{i,T}(t)/\gamma_{T}-(1/\gamma_{T})\nabla_{\mathbf{r}_{i}}U,$ (1)
@@
where $\gamma_{T},\gamma_{R}$ are the Stokes friction coefficients associated with translation (T) and rotation (R) in the $x-y$ plane, respectively, $u_{0}$ represents the self-propelled velocity with constant magnitude see Figure 1. Also, in Eqs.(1) and (2), $\bm{\xi}_{T}(t)$ and $\bm{\xi}_{R}(t)$ are, respectively, the translational and rotational, Gaussian, white noises with zero mean, $\langle\bm{\xi}_{T}(t)\rangle=\langle\bm{\xi}_{R}(t)\rangle=\mathbf{0}$, and time correlation $\langle\bm{\xi}_{T}(t_{1})\bm{\xi}_{T}(t_{2})\rangle=D_{T}\delta(t_{1}-t_{2})$ and $\langle\bm{\xi}_{R}(t_{1})\bm{\xi}_{R}(t_{2})\rangle=D_{R}\delta(t_{1}-t_{2})$, where we can define $D_{T}$ and $D_{R}$ as the translation and rotation diffusion parameters. Finally, in Eq.(1), $U$ corresponds to the summation over steric interactions between particles at time $t$,
@@
with $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ is the distance between particles, $\sigma$ the particle diameter, $C_{r}=2^{1/6}\sigma\approx 1.1225\sigma$ is defined here as the contact radius, and $\varepsilon$ is the energy interaction strength between particles. We set $\varepsilon=3$ during our simulations unless we stated otherwise.
Figure 1: Time evolution for a system of active Brownian particles exhibiting Motility-induced phase separation: In the top-left, a single ABP is shown with orientation vector $\hat{u}_{i}$ and radius $R$. The time evolution is shown at three different times, showing a transition from a single phase in $t=0.02t_{r}$ and two coexisting phases in $t=5.85t_{r}$, at $\text{Pe}=140$ and $\phi=0.5$, with $L=100$ and $t_{r}=1/D_{R}$.
@@
The interplay between steric interactions causing the blocking of persistent motion and an increased probability of collision due to elevated concentrations $\phi$ and particle activities $u_{0}$ leads to a fascinating phenomenon known as motility-induced phase separation (MIPS) theers2018clustering ; stenhammar2014phase . The system starts with a homogeneous particle distribution and orientations at $t=0$ exhibiting a single-phase behavior (or fluid-like state su2023motility ), which physically describes ABP randomly moving without particle agglomeration. However, if the packing fraction and Péclet number are larger than some critical values, the system evolves into a two-phase state represented by a solid and gas phase, as illustrated in Figure 1. In this phase, particles agglomerate, and cluster formation becomes an emergent phenomenon.
@@
In Fig. 2, we analytically reproduced the phase diagram in terms of the packing fraction and the Péclet number using $\kappa=4.7$ as a fit parameter redner2013structure . The pairs $(\text{Pe},\phi)$ studied in this work are represented by the white dots of Fig. 2. For $f_{c}=0$, we can say that the system is in the single-phase, while for $f_{c}\neq 0$, the system undergoes a phase transition to the two-phase state. The system is considered steady when the phase separation shows a bimodal probability distribution for different local packing fractions redner2013structure ; stenhammar2014phase . Despite extensive studies on phase separation in various contexts, machine learning has recently investigated the time evolution during phase separation, revealing different transition regimes mcdermott2023characterizing . In this work, we aim to study the time evolution by generating a temporal averaged network during the phase transition.
Figure 2: An analytical representation for the phase diagram in the $\text{Pe}-\phi$ plane plotted from the analytical model using Eq. 5, with $\kappa=4.7$. The white dots represent the pairs $(\text{Pe},\phi)$ parameters studied across this work.
2.1 Simulation methods and parameters
@@
The relevant length scale in the system is the particle radius $R$, serving as the unit of length. Additionally, we adopted $t_{r}$ as the unit of time. The equations of motion, Eqs.(1) and (2), were integrated with a time step of $\Delta t=3\times 10^{-5}t_{r}$, and simulation times were set to $T=12t_{r}$. Two dimensionless values parameterize the system: the packing fraction $\phi$ and the Péclet number, which are varied in the intervals $\phi\in[0.3-0.7]$ and $\text{Pe}\in[80,140]$, respectively (see Fig. 2). We maintain a constant rotational diffusion coefficient $D_{R}=0.15$, translation diffusion coefficient $D_{T}=0.2$, and diameter $\sigma=2R$ with $R=1$.
3 Complex Networks for ABP
A network is defined as a group of nodes connected between them by edges. The connections between nodes could be directed or undirected. In this case, the nodes are determined by each ABP, and the interaction between particles defines the connections: two nodes (particles) are connected when their centers are inside the interaction region defined by the contact radius $C_{r}$ (see WCA potential in Eq. (4)). We build a time network by following the dynamics of the particle connections in a specific time window.
@@ 3 Complex Networks for ABP
Figure 3: From left to right: A scheme for particle encounters when a pair interaction occurs. The blurred area represents the interaction region with radius $C_{r}$. Particles (nodes) generate a link if their centers enter each other’s interaction region, as shown in the middle panel. A representation of the time evolution for the network formation, with time windows of length $W=0.45t_{r}$ and overlap $t_{0}=0.225t_{r}$ between two consecutive windows.
3.1 Adjacency matrix
In order to analyze the system, we first create the adjacency matrix $A$, which is used to represent the network. Using the distance $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ between the centers of two particles, we can establish the value of $A_{ij}$ which represents the connections between these particles. If $r_{ij}\leq C_{r}$, we set $A_{ij}=A_{ji}=1$; otherwise, $A_{ij}=0$. If $A_{ij}=A_{ji}=1$ it means that two nodes (particles) are connected between them, and the complex network built is undirected. Within time windows of length $W=0.45t_{r}$ and using an overlapping time $t_{0}=0.225t_{r}$, a link is established only upon the first interaction between two particles. The latter means that each network accounts for unique particle encounters within the time window. From the adjacency matrix, we can compute all the necessary properties of the network. We will focus on calculating the degree distribution, network average clustering coefficient, and average path length.
3.2 Degree distribution
The degree $deg(p_{i})$ of each particle $p_{i}$ is the number of connections of a node, and it can be computed by summing over the corresponding column or row of the adjacency matrix:
@@ 3.2 Degree distribution
which is directly obtained from simulations. Here $n_{k}$ is the total number of particles with degree $k$, and $N$ is the number of particles in the system. Depending on the topology of the complex network, the degree distribution could be binomial (random graphs), power-law (scale-free), among others. In this work, we shall discuss the role of the Gaussian distribution and double power law models in the context of MIPS.
3.3 Network average clustering coefficient
The clustering coefficient $C_{i}$ of each particle is a measure of the tendency to form clusters of this node, and it is defined by
@@ 3.3 Network average clustering coefficient
where $A$ is the adjacency matrix, and $k_{i}$ is the degree of the particle. From this, we can calculate the network average clustering coefficient $\bar{C}$, which is the trend of the network to form clusters, and it is defined by
@@
3.4 Average path length
The average path length is the rate between the shortest number of steps between two nodes and the shortest paths between all the pairs of nodes, and it is defined by Newman_2000 ,
@@ 3.4 Average path length
Here, $d(v_{i},v_{j})$ is the shortest distance between two vertices $v_{i}$ and $v_{j}$, where $v_{i,j}$ are nodes of the graph.
4 Results and Discussion
We studied three different regimes of the ABP according to the phase diagram of Fig. 2: (i) one-phase (violet), (ii) transient phase (blue-green), and (iii) two-phase (orange-red). As discussed in the following sections, these three regions will exhibit different topological properties, which are not reported in the literature to the best of our knowledge.
4.1 One-phase region: Random Graph
@@
where $\sigma_{\rm st}$ is the standard deviation, and $\mu$ is the average number of unique encounters of particles in a time window $W$. We have a dilute system for packing fraction numbers $\phi<0.4$, which allows us to analyze the average behavior of the ABP system using the kinetic theory of gases. As explained in Ref. norambuena2020understanding , in an active media with $N$ moving particles in a dilute system, the mean free path is $\lambda=\sqrt{2}u_{0}\tau_{c}$, where $\tau_{c}$ is the mean time between encounters. During a time interval $\tau_{c}$, we can establish the probability constraint $p_{c}=A_{\rm sw}/L^{2}\approx 1$, where $A_{\rm sw}=N(2\lambda a+\pi a^{2})$ is the total area swept area of the $N$ particles, and $L^{2}$ is the available area. From this argument, it follows that $\lambda=(\pi/4)\phi^{-1}(1-\phi)\sigma=\sqrt{2}u_{0}\tau_{c}$. Then, the average number of encounters can be estimated as $\mu=W/\tau_{c}$, since, on average (by definition), particles have one encounter in the time interval $\tau_{c}$. We found:
$\mu(\phi,\text{Pe})={4\sqrt{2}\over\pi}\left({\phi\over 1-\phi}\right)WD_{R}\text{Pe},$ (12)
@@
Figure 4: Degree distribution $P(k)$ for systems in the one-phase region at $\text{Pe}=80$ (circles) and $\text{Pe}=140$ (diamonds) at time $t=0.45t_{r}$ (top) and at $t=11.0t_{r}$ (bottom). The distributions show a Gaussian tendency; each system exhibits distinct mean and standard deviation values. Note that all systems maintain a Gaussian distribution, and only the system at $\phi=0.4$ shows a change over time. Both insets (top and bottom) show the $\log-\log$ plot of the systems at the respective time.
@@
where $W=0.45t_{r}$ is fixed for $D_{R}=0.15$. In Fig. 4, we illustrate the evolution of the degree distribution $P(k)$ measured from Eqs. 6 and 7 for a system at two time lapses, $t=0.45t_{r}$ and $t=11.0t_{r}$. Two Péclet numbers are considered: $\text{Pe}=80$ represented by circles and $\text{Pe}=140$ represented by diamonds. Different packing fractions are depicted with orange and blue colors, corresponding to $\phi=0.3$ and $\phi=0.4$, respectively. Most importantly, using our model given in Eq. 12, we obtain $\mu(\phi=0.3,\text{Pe}=80)\approx 10.3$ and $\mu(\phi=0.3,\text{Pe}=140)\approx 16.2$, whose numerical values are close to those obtained by simulations ($10.51$ and $16.75$), see Fig. 5. This simple model can give good estimations of the central value of the Gaussian distribution $P^{\rm G}(k)$ given in equation 11, which is valid for the one-phase region.
Figure 5: Standard deviation $\sigma_{st}$ (top) and mean $\mu$ (bottom) values over time of Gaussian fits for the degree distribution of systems in the one-phase region (Fig. 4). Both $\sigma_{st}$ and $\mu$ remain virtually constant over time and only the mean for the system at $\phi=0.4$ shows an appreciable decrease.
4.2 Transient Region: Hybrid phase

### v2

# Emerging double power-law through dynamical complex networks in motility-induced phase separation in Active Brownian Particles

@@ Abstract
We investigate the behavior of active Brownian particles (ABP) within a temporal complex network framework approach. We focused on the node degree distribution, average path length, and average clustering coefficient across the Péclet number and packing fraction region. In the single phase or gas region, particle interactions mirror a random graph, and the average number of unique interactions display a non-monotonic behavior with the packing fraction. As we ventured toward the Motility-Phase Separation (MIPS) frontier, a hybrid distribution with a time-evolving pattern emerged, combining Gaussian and power law components. Moreover, we discovered a double power-law distribution in the phase-separated region, with two characteristic slopes symbolizing the emerging gas and solid regions. Our approach involved various numerical and theoretical analyses to capture the role of the packing fraction and Péclet number in the topological properties of the dynamical complex network that arises during the ABP dynamics. These findings provide a robust understanding of a phase transition between two states and how this transition manifests as a random to scale-free behavior.
preprint: AIP/123-QED
I Introduction
Active Brownian Particles (ABP) serve as a valuable framework for investigating the non-equilibrium dynamics and statistical properties of self-propelled agents, whether they are living organisms or artificial entities romanczuk2012active ; bechinger2016active . In contrast to passive Brownian particles studied in random diffusive systems, ABP offers a new dimension of complexity due to the inherent particle self-propulsion mechanisms. In particular, these active particles undergo a distinctive phase transition known as Motility-Induced Phase Separation (MIPS) fily2012athermal ; redner2013structure ; stenhammar2014phase ; theers2018clustering ; van2019interrupted . This phase transition is related to the spontaneous separation between dilute (free-particles) and dense phases (cluster formation) stenhammar2014phase ; de2023sequential ; de2021active . This phase separation is a fascinating emergent phenomenon where particles collapse into clusters although natural repulsion forces are present. The exploration of ABP not only deepens into the understanding of self-organization and emergent phenomena but also opens up promising avenues for harnessing their unique properties in various applications and technological advancements gompper20202020 and building a bridge between physical models and biological systems.
@@ I Introduction
II Active Brownian Particles (ABP) and motility-phase-separation
@@
$\dot{\mathbf{r}}_{i}$ $=$ $u_{0}\hat{\mathbf{u}}_{i}(t)+(1/\gamma_{T})(\bm{\xi}_{i,T}(t)-\nabla_{\mathbf{r}_{i}}U),$ (1)
@@
where $\gamma_{T},\gamma_{R}$ are the Stokes friction coefficients associated with translation (T) and rotation (R) in the $x-y$ plane, respectively, $u_{0}$ represents the self-propelled velocity with constant magnitude see Fig. 1. Also, in Eqs.(1) and (2), $\bm{\xi}_{T}(t)$ and $\bm{\xi}_{R}(t)$ are, respectively, the translational and rotational, Gaussian, white noises with zero mean, $\langle\bm{\xi}_{T}(t)\rangle=\langle\bm{\xi}_{R}(t)\rangle=\mathbf{0}$, and time correlation $\langle\bm{\xi}_{T}(t_{1})\bm{\xi}_{T}(t_{2})\rangle=D_{T}\delta(t_{1}-t_{2})$ and $\langle\bm{\xi}_{R}(t_{1})\bm{\xi}_{R}(t_{2})\rangle=D_{R}\delta(t_{1}-t_{2})$, where we can define $D_{T}$ and $D_{R}$ as the translation and rotation diffusion parameters. Finally, in Eq.(1), $U$ corresponds to the summation over steric interactions between particles at time $t$,
@@
with $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ is the distance between particles, $\sigma$ the particle diameter, $C_{r}=2^{1/6}\sigma\approx 1.1225\sigma$ is defined here as the contact radius, and $\varepsilon$ is the energy interaction strength between particles. We set $\varepsilon=3$ unless we stated otherwise during our simulations.
Figure 1: Time evolution for a system of active Brownian particles exhibiting Motility-induced phase separation: In the top-left inset, a single ABP is shown with orientation vector $\hat{u}_{i}$ and radius $R$. The time evolution is shown at four different times, showing a transition from a single phase in $t=0.0t_{r}$ and two coexisting phases in $t=6.0t_{r}$, at $\text{Pe}=140$ and $\phi=0.5$, with $L=100$ and $t_{r}=1/D_{R}$.
@@
The interplay between steric interactions causing the blocking of persistent motion and an increased probability of collision due to elevated concentrations $\phi$ and particle activities $u_{0}$ leads to a fascinating phenomenon known as motility-induced phase separation (MIPS) theers2018clustering ; stenhammar2014phase . The system starts with a homogeneous particle distribution and orientations at $t=0$ exhibiting a single-phase behavior (or fluid-like state su2023motility ), which physically describes ABP randomly moving without particle agglomeration. However, if the packing fraction and Péclet number are larger than some critical values, the system evolves into a two-phase state represented by a solid and gas phase, as illustrated in Fig 1. In this phase, particles agglomerate, and cluster formation becomes an emergent phenomenon.
@@
Figure 2: Analytical representation for the phase diagram in the $\text{Pe}-\phi$ plane plotted from the analytical model using Eq. (5), with $\kappa=4.7$ redner2013structure . The white dots represent the pairs $(\text{Pe},\phi)$ parameters studied across this work. The different colors represent the observed networks. White dots represent the single phase where a random graph is observed. The yellow dots represent the hybrid phase, and the red dots represent the two-phase region where a double power-law distribution is recovered.
In Fig. 2, we analytically reproduced the phase diagram in terms of the packing fraction and the Péclet number using $\kappa=4.7$ as a fit parameter redner2013structure . The pairs $(\text{Pe},\phi)$ studied in this work are represented by dots in Fig. 2. For $f_{c}=0$, we can say that the system is in the single-phase (white dots), while for $f_{c}\neq 0$, the system undergoes a phase transition to the two-phase state (red dots). The system is considered steady when the phase separation shows a bimodal probability distribution for different local packing fractions redner2013structure ; stenhammar2014phase . Despite extensive studies on phase separation in various contexts, machine learning has recently investigated the time evolution during phase separation, revealing different transition regimes mcdermott2023characterizing . In this work, we aim to study the time evolution by generating a temporal averaged network during the phase transition.
II.1 Simulation methods and parameters
@@
The relevant length scale in the system is the particle radius $R$, serving as the unit of length. Additionally, we adopted $t_{r}$ as the unit of time. The equations of motion, Eqs.(1) and (2), were integrated with a time step of $\Delta t=3\times 10^{-5}t_{r}$, and simulation times were set to $T=30t_{r}$. Two dimensionless values parameterize the system: the packing fraction $\phi$ and the Péclet number, which are varied in the intervals $\phi\in[0.3-0.7]$ and $\text{Pe}\in[80,140]$, respectively (see Fig. 2). We maintain a constant rotational diffusion coefficient $D_{R}=0.15$, translation diffusion coefficient $D_{T}=0.2$, the diameter $\sigma=2R$ with $R=1$, and the Stokes translational and rotational friction coefficients are $\gamma_{T}=6\pi$ and $\gamma_{R}=8\pi$ respectively.
III Complex Networks for ABP
A network is defined as a group of nodes connected between them by edges. The connections between nodes could be directed or undirected. In this case, the nodes are determined by each ABP, and the interaction between particles defines the connections: two nodes (particles) are connected when their centers are inside the interaction region defined by the contact radius $C_{r}$ (see WCA potential in Eq. (4)). We build a time network by following the dynamics of the particle connections in a specific time window. The network is built considering a fixed number of nodes $N$ and an increasing number of connections.
@@ III Complex Networks for ABP
III.1 Adjacency matrix
In order to analyze the system, we first create the adjacency matrix $A$, which is used to represent the network. Using the distance $r_{ij}=|\mathbf{r}_{i}-\mathbf{r}_{j}|$ between the centers of two particles, we can establish the value of $A_{ij}$ which represents the connections between these particles. If $r_{ij}\leq C_{r}$, we set $A_{ij}=A_{ji}=1$; otherwise, $A_{ij}=0$. If $A_{ij}=A_{ji}=1$ it means that two nodes (particles) are connected between them, and the complex network built is undirected. Within time windows of length $W=0.45t_{r}$ and using an overlapping time $t_{0}=0.225t_{r}$, a link is established only upon the first interaction between two particles. The latter means that each network accounts for unique particle encounters within the time window.
Figure 3: From left to right: Scheme for particle encounters when a pair interaction occurs. The blurred area represents the interaction region with radius $C_{r}$. Particles (nodes) generate a link if their centers enter each other’s interaction region, as shown in the middle panel. A representation of the time evolution for the network formation, with time windows of length $W=0.45t_{r}$ and overlap $t_{0}=0.225t_{r}$ between two consecutive windows.
From the adjacency matrix, we can compute all the necessary properties of the network. As we know, the study of complex networks can shed light on the global behavior of a system composed of small and separated elements. We will focus on calculating the degree distribution, network average clustering coefficient, and average path length. These metrics give good insights into describing, characterizing, and understanding the time evolution of the structure of a complex network.
III.2 Degree distribution
As a first manner to characterize these complex networks, we compute the degree of the nodes. We define the degree $deg(p_{i})$ of each particle $p_{i}$ as the number of connections of a node, and it can be computed by summing over the corresponding column or row of the adjacency matrix:
@@ III.2 Degree distribution
which is directly obtained from simulations. Here $n_{k}$ is the total number of particles with degree $k$, and $N$ is the number of particles in the system. Depending on the topology of the complex network, the degree distribution could show different behaviors Erdos ; barabasi1999 ; albert2002statistical ; newman2003 : it could be binomial (random graphs) or power-law (scale-free). Real networks usually show a scale-free behavior. To explain this behavior of a complex network built with the World Wide Web, Réka Albert and Albert Barabási barabasi1999 ; albert2002statistical ; barabasi2011 used a model for random networks given a preferential probability of attachment for the new connections. In this work, we shall discuss the role of the Gaussian distribution and double power law models in the context of MIPS to understand the global structure of these complex networks. Also, the degree distribution has been shown to be useful to identify phase transitions in real networks pasten2016 .
III.3 Network average clustering coefficient
Another characteristic of real networks is the small-world phenomenon. Watts and Strogatz watts1998 show that random networks do not have this behavior in which two nodes of the complex network are connected between them through small steps. In order to give a measure of the small world behavior, we can define a global clustering coefficient (for all the nodes of the network) and a local clustering coefficient (for each node). The values of the clustering coefficient allow us to identify these small-world properties. Both random networks and ordered lattices do not show a small-world behavior. The clustering coefficient Newman_2000 ; newman2003 $C_{i}$ of each particle is a measure of the tendency to form clusters of this node, and it is defined by
@@ III.3 Network average clustering coefficient
where $A$ is the adjacency matrix, and $k_{i}$ is the degree of the particle. From this, we can calculate the network average clustering coefficient watts1998 $\bar{C}$, which is the trend of the network to form clusters, and it is defined by
@@
III.4 Average path length
To identify the small-world phenomenon, we need an additional measure: the distance between two nodes, called average path length Newman_2000 . The small-world model shows an increase in the log scale of the average path length and the clustering properties at the same time. The average path length is the rate between the shortest number of steps between two nodes and the shortest paths between all the pairs of nodes, and it is defined by Newman_2000 ,
@@ III.4 Average path length
Here, $d(v_{i},v_{j})$ is the shortest distance between two vertices $v_{i}$ and $v_{j}$, where $v_{i,j}$ are nodes of the graph. The small-world model shows a single length scale that depends on the probability of the connections between nodes newman_watts1999 .
IV Results and Discussion
We studied three different regimes of the ABP according to the phase diagram of Fig. 2: (i) one-phase (white circles), (ii) transient phase (yellow circles), and (iii) two-phase (red circles). As discussed in the following sections, these three regions will exhibit different topological properties.
IV.1 One-phase region: Random Graph
@@
[…]
$\mu(\phi,\text{Pe})=\Lambda(\text{Pe})\left({\phi\over 1-\phi}\right)(1-\chi\phi)P(\lambda),$ (12)
@@
Figure 4: (a), (b) Time evolution for the degree distribution $P(k)$ for systems in the one-phase region at $\text{Pe}=20$ (triangles) and $\text{Pe}=140$ (circles) at time $t=0.45t_{r}$ and at $t=30.0t_{r}$ for different packing fractions $\phi$. The distributions show a Gaussian tendency; each system exhibits distinct mean and standard deviation values. Note that all systems maintain a Gaussian distribution. The inset shows the average number of unique encounters between particles $\mu$, for different Pe and $\phi$, extracted from (c). (c) The average number of unique encounters of particles $\mu$ for different Péclet numbers in the single-phase, dashed lines show the analytical value following Eq. (12).
@@
where $\Lambda(\text{Pe})=4\sqrt{2}WD_{R}\text{Pe}/\pi$, $P(\lambda)=1-e^{-(\lambda-\sigma)/\lambda}$, $\lambda=\lambda(\phi)$, $W=0.45t_{r}$, $D_{R}=0.15$, and we use $\chi=0.2$ to have a better fit. In Fig. 4 (a), (b), we show the time evolution for the Gaussian distributions of $P(k)$ for different packing fractions and Péclet numbers. Using our model given in Eq. (12), we obtain $\mu(\phi=0.1,\text{Pe}=140)\approx 7.118$, $\mu(\phi=0.2,\text{Pe}=140)\approx 13.456$, and $\mu(\phi=0.3,\text{Pe}=140)\approx 16.686$ whose values are close to those obtained by our simulations ($6.804$, $12.659$, $16.773$), see Fig. 4 (c). This simple model can give good estimations of the central value of the Gaussian distribution $P^{\rm G}(k)$ given in Eq. (11) in the regime $\phi<\phi_{c}=0.41$. The inset in Fig. 4 (b) resumes the obtained values for $\mu$ in our simulations. Showing a non-monotonic behavior, where $\mu$ increases with $\phi$ for dilute concentrations and $\mu$ decreases for denser systems.
IV.2 Transient Region: Hybrid phase